OpenStudy (bahrom7893):

Kelsey, continue here: An auditorium earned $30,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (bahrom7893):

75f+50b = 30,000 f=2b

OpenStudy (bahrom7893):

b = f/2 75f + 25f = 30,000 100f = 30,000 f = 300

OpenStudy (bahrom7893):

300s the answer

OpenStudy (anonymous):

thankyou so much, i really appreciate everything.

OpenStudy (anonymous):

I LOVE YOU!...now to submit my test.

OpenStudy (bahrom7893):

drumroll...

OpenStudy (anonymous):

lol